OpenStudy (briana.img):

Kevin has 23 dimes and quarters in his piggy bank. If the value of these coins is $4.55, how many dimes and quarters does he have? Set up and solve a system of equations to solve the problem.

OpenStudy (briana.img):

Been doing this problem for a while but I always get stuck at the distributive property

OpenStudy (anonymous):

that is a tough property that one how far did you get?

OpenStudy (briana.img):

x+y=42 7x+3y=194 7(42-y)+3y=194

OpenStudy (anonymous):

where in the world did \(7x+3y=194\) come from ?

OpenStudy (briana.img):

oh wait this is my work for another problem sorry

OpenStudy (anonymous):

\[x+y=23\\ 25x+10y=455\]

OpenStudy (briana.img):

substitution method?

OpenStudy (anonymous):

sure i would have gone right to \[25x+10(23-x)=455\] and skipped the first silliness why use two variables when one will do

OpenStudy (briana.img):

25(23-y)+10y=455

OpenStudy (briana.img):

oh you're doing x first

OpenStudy (anonymous):

whatever i solve for \(x\) you solve for \(y\) potato potahto

OpenStudy (briana.img):

25(23-y)+10y=455 575-23y=455 i always get messed up at this part

OpenStudy (anonymous):

whoa nellie hold the phone we are doing yours right ?\[25(23-y)+10y=455 \]

OpenStudy (anonymous):

leave the \(10y\) term alone for a moment and multiply \[25(23-y)\]

OpenStudy (briana.img):

575y ?!

OpenStudy (anonymous):

yeah i see that also you have \(-23y\) which is also wrong

OpenStudy (briana.img):

yeah i haven't understood that fully

OpenStudy (anonymous):

\[25(23-y)=25\times 23-25\times y\]

OpenStudy (anonymous):

\[a(b-c)=ab-ac\] that is the distributive law

OpenStudy (briana.img):

oh ok

OpenStudy (anonymous):

\[\huge \heartsuit (\clubsuit+\diamondsuit)=\heartsuit\times \clubsuit+\heartsuit\times \diamondsuit\]

OpenStudy (anonymous):

so the first step should be \[25(23-y)+10y=455\\ 25\times 23-23y+10y=455\] then compute and combine like terms

OpenStudy (briana.img):

575-33y=455

OpenStudy (anonymous):

let me know when you get \[575 - 13 y == 455\]

OpenStudy (radar):

typo change -23y to read -25y

OpenStudy (anonymous):

oh a see another problem here

OpenStudy (anonymous):

yeah i made a typo should be \[25(23-y)+10y=455\\ 25\times 23-25y+10y=455\]

OpenStudy (anonymous):

then \[575 - 15 y = 455\]

OpenStudy (briana.img):

then subtract 575 from 455?!

OpenStudy (anonymous):

note that \(-25y+10y=-15y\) it is the same as \(10y-25y\)

OpenStudy (anonymous):

yes

OpenStudy (briana.img):

15y=120 y=8

OpenStudy (briana.img):

x+8=23 x=15

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i mean yes you are right